Given : 1/16 ÷ 1/81 + - 1/8
 
To find: multiplicative inverse

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
a multiplicative inverse is basically a reciprocal

Multiplicative inverse of number is the number which if multiplied by original number result in 1

1/16 ÷ 1/81 + -1/8

1/16 ÷ 1/81 = 81/16

+ -1/8 = - 1/8

81/16 - 1/8

= 81/16 - 2/16

= (81 - 2)/16

= 79/16

Multiplicative inverse = 16/79